a train covers a distance of 220 mi in 2hr and 45min.calculate the average speed of the train in m/s.
Answer by mananth(16949)

d=220
t= 2 45 minutes = 2 3/4 hours=> 11/4 hours
r=d/t
r=220/(11/4)
r=220*4/11
r=80 mph
1 mile = 1600 m
80 miles = 128000 m
1 hour = 3600 seconds
r=128000/3600
r=35.55m/s